# Density control of multi-agent swarms via bio-inspired leader-follower plasticity

### v1

@@ 6 Higher-Dimensional Extension
Next, we extend the theoretical framework to periodic domains in higher-dimensions, $\Omega=[-\pi,\pi]^{d}$ ($d=2,3$).
The model in (1) becomes
$\rho^{L}_{t}(\mathbf{x},t)+\nabla\cdot[\rho^{L}(\mathbf{x},t)$ $\mathbf{u}(\mathbf{x},t)+\rho^{L}(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)]=$
$=D\nabla^{2}\rho^{L}(\mathbf{x},t)+q(\mathbf{x},t),$ (58a)
$\rho^{F}_{t}(\mathbf{x},t)+\nabla\cdot[\rho^{F}(\mathbf{x},t)$ $(\mathbf{f}*\rho)(\mathbf{x},t)]=D\nabla^{2}\rho^{F}(\mathbf{x},t)$
$\qquad\qquad-q(\mathbf{x},t),$ (58b)
$\eta^{F}_{t}(\mathbf{x},t)+\nabla\cdot[\eta^{F}(\mathbf{x},t)$ $(\mathbf{f}*\rho)(\mathbf{x},t)]=D\nabla^{2}\eta^{F}(\mathbf{x},t),$ (58c)
@@ 6.1 Bio-inspired Control
$\nabla\cdot\left[\rho^{L}(\mathbf{x},t)\mathbf{u}(\mathbf{x},t)\right]=-Ke(\mathbf{x},t)\\ -\nabla\cdot\left[\rho(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)\right]-D\nabla^{2}\rho(\mathbf{x},t)$ (59)
@@
The error dynamics obeys the higher-dimensional extension of (11). Choosing (59) brings the error dynamics in the form of (14), proving the claim. ∎
@@
To uniquely recover $\mathbf{u}$, a vector field, from the scalar relation (59), extra constraints need to be included, simlar to [44]. Specifically, we set $\mathbf{w}:=\rho^{L}\mathbf{u}$, and
$Y(\mathbf{x},t)=-Ke(\mathbf{x},t)-\nabla\cdot\left[\rho(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)\right]\\ -D\nabla^{2}\rho(\mathbf{x},t),$ (60)
@@
$\begin{cases}\nabla\cdot\mathbf{w}(\mathbf{x},t)=Y(\mathbf{x},t),\\ \nabla\times\mathbf{w}(\mathbf{x},t)=0,\end{cases}$ (61)
where we added a zero-curl condition to (59). Such a problem is analogous to the Poisson equation $\nabla^{2}\varphi=-Y$, where $\mathbf{w}=-\nabla\varphi$. The Poisson problem can be solved using Fourier series expansion, so that $\varphi$, and consequently $\mathbf{w}$ are recovered, from which one establishes $\mathbf{u}=\mathbf{w}/\rho^{L}$.
@@
$q(\mathbf{x},t)=\frac{1}{2}\,\nabla\cdot\left[\rho^{L}(\mathbf{x},t)(\mathbf{x},t)\right]+\frac{1}{2}\,\nabla\cdot\left[\rho^{*}(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)\right]\\ -\frac{D}{2}\nabla^{2}\rho^{*}(\mathbf{x},t)+g(\mathbf{x},t),$ (62)
@@ 6.2 Stability Analysis
$\int(f_{1}*\psi)(x_{1},x_{2},x_{3})\,\mathrm{d}x_{1}=\int(f_{2}*\psi)(x_{1},x_{2}x_{3})\,\mathrm{d}x_{2}=\\ =\int(f_{3}*\psi)(x_{1},x_{2},x_{3})\,\mathrm{d}x_{3},$ (63)
for any periodic $\psi$. Choosing $\mathbf{u}$ according to (59) (see Rem. 10) and $q$ as in (62) implies that $\bar{\rho}$ is a steady-state solution for the dynamics of $\rho$, with $\rho^{L},\rho^{F}>0$, and $\eta^{F}\geq 0$, if and only if the higher-dimensional extension of (20) holds.
@@
Under the additional assumption of isotropic interaction kernel, the proof follows the same steps of those in Theorem 4. The only difference is the computation of the steady-state solution of $\eta^{F}$ (see (27)). Setting $\eta^{F}_{t}=0$ and $\rho=\bar{\rho}$ in (58c), we obtain
$\nabla\cdot\left[\bar{\eta}^{F}(\mathbf{x})(\mathbf{f}*\bar{\rho})(\mathbf{x})\right]$ $=D\nabla^{2}\bar{\eta}^{F}(\mathbf{x}),$ (64)
@@
$\nabla\cdot\left[\bar{\eta}^{F}(\mathbf{x})(\mathbf{f}*\bar{\rho})(\mathbf{x})-D\nabla\bar{\eta}^{F}(\mathbf{x})\right]=0.$ (65)
Equation (65) is fulfilled if [^2] This condition becomes also necessary when $\bar{\eta}^{F}$, $\mathbf{f}$ and $\bar{\rho}$ satisfy ×$\nabla\times\left[\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}\right]=0$. This is equivalent to requiring that $\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}=-\nabla\Psi$, where $\Psi$ is a harmonic scalar potential. Under our isotropic kernel assumption, the solution we construct satisfies this condition, ensuring the uniqueness of the steady-state solution (see the following analysis and Remark 12).
$\nabla\bar{\eta}^{F}(\mathbf{x})=\frac{1}{D}\bar{\eta}^{F}(\mathbf{x})(\mathbf{f}*\bar{\rho})(\mathbf{x}).$ (66)
@@
Many interaction kernels from the literature satisfy condition (63). Also, under this hypothesis, the steady-state solution for $\bar{\eta}^{F}$ is uniquely defined in higher dimensions, since $\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}=-\nabla\Psi$, where $\Psi$ is a harmonic scalar potential.
@@
the error dynamics locally converges to 0 in $\mathcal{L}^{2}(\Omega)$.
@@
The proof follows the same structure of that of Theorem 8. Dropping time and space dependencies for simplicity, the time derivative of the Lyapunov functional in (44) can be rewritten as
$V_{t}=-2D\|\nabla e^{\eta}\|_{2}^{2}-\int_{\Omega}\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\,\mathrm{d}\mathbf{x}\\ -2\mathrm{exp}(-Kt)\int_{\Omega}e^{\eta}\tilde{\mathbf{v}}\,\mathrm{d}\mathbf{x},$ (71)
using the divergence theorem, vectorial identities, and posing $\tilde{\mathbf{v}}=\nabla\cdot(\bar{\eta}^{F}\,\mathbf{f}*e^{0})$. The first term on the right-hand side can be bounded using Poincaré-Wirtinger inequality, and for the second one the following bound holds:
$\left|\int_{\Omega}\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\,\mathrm{d}\mathbf{x}\right|\leq\int_{\Omega}\left|\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\right|\,\mathrm{d}\mathbf{x}=\\ =\left\|e^{\eta}e^{\eta}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\right\|_{1}\leq\|e^{\eta}\|_{2}^{2}\|\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\|_{\infty}\leq\\ \leq\|e^{\eta}\|_{2}^{2}\sum_{i=1}^{d}\|(f_{i}*\bar{\rho}_{x_{i}})\|_{\infty}\leq\|e^{\eta}\|_{2}^{2}\sum_{i=1}^{d}\|f_{i}\|_{2}\|\bar{\rho}_{x_{i}}\|_{2},$ (72)
@@
$\left|-2\mathrm{exp}(-Kt)\int_{\Omega}e^{\eta}\tilde{\mathbf{v}}\,\mathrm{d}\mathbf{x}\right|\leq 2\mathrm{exp}(-Kt)\|e^{\eta}\tilde{\mathbf{v}}\|_{1}\leq\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}\|_{2}\|\tilde{\mathbf{v}}\|_{2}$ (73)
@@
$V_{t}\leq\left(-2D+\sum_{i=1}^{d}\|f_{i}\|_{2}\|\bar{\rho}_{x_{i}}\|_{2}\right)V\\ +2\mathrm{exp}(-Kt)\|\tilde{\mathbf{v}}\|_{2}\sqrt{V}.\qed$ (74)
@@ 6.3 Numerical validation
We consider a 2D mono-modal regulation scenario with a bivariate von Mises distribution (zero means, unit concentration). Agents interact through a 2D periodic Morse kernel ($L_{a}=\pi$, $L_{r}=\pi/4$, $\alpha=3.2$) with $D=0.05$, $\Phi^{F}=0.2$, $M^{L}(0)=M^{F}(0)=0.4$, $K=1$, $K_{FL}=1$, $K_{LF}=2$. Results in Fig. 4 for $t_{\mathrm{f}}=100$ are qualitatively comparable to those of 1D simulations in Figs 1. Convergence of $\rho$ to $\bar{\rho}$ occurs in  5 time units while convergence of $\rho^{L}$, $\rho^{F}$ and $\eta^{F}$ to their steady-state profiles is slower. We obtain final masses $M^{L}(t_{\mathrm{f}})\approx 0.26$ and $M^{F}(t_{\mathrm{f}})\approx 0.53$, consistent with the predicted $1/2$ ratio of in Corollary 6.
@@ 7 Conclusions
We presented a novel bio-inspired leader-follower technique to control the spatial organization of large swarms of mobile agents. Upon formulating a density control problem, we derived conditions for the existence and local stability of desired solutions. Our strategy incorporates two crucial, yet previously overlooked, characteristics of natural systems. Namely, we set control objectives for the entire collective, including both leaders and followers, and we introduce behavioral plasticity to allow for tuning the leaders-to-follower mass ratio. We provide necessary and sufficient conditions for the existence of solutions to our control problem and assess their local stability, both for 1D and higher-dimensional settings. Our numerical findings offer compelling support to the mathematical derivations and to the effectiveness, robustness, and versatility of the proposed bio-inspired control strategy.
Our work does not come without limitations. Our analytical guarantees of convergence hold at the scale of PDEs, that is, in the limiting scenario of swarms of infinite size. The convergence from the continuum macroscopic PDE model to the discrete microscopic agent-based implementation should be assessed; potentially, this may be tackled through two-scale [45] and $\Gamma$-convergence [46]. Moreover, our theoretical results yield exact predictions only for steady-state leaders-to-followers mass ratio. Although numerical simulations suggest a monotonic trend in the leaders’ and plastic followers’ masses, we cannot argue that these trends are universal. As such, we may not exclude cases in which there is considerably higher role-switching during the system evolution than steady-state predictions. Exploring these transient dynamics may be valuable in the characterization of energetic costs of behavioral plasticity. Beyond addressing these limitations, we envision future work in which continuum descriptions of agents are not solely used for spatial organization, but to accomplish more complex tasks, such as self-assembly, collaborative manipulation, and object clustering [47].
@@ Appendix A Non-plastic followers at steady-state
$\bar{\eta}^{F}(x)=B\,\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})(x)\,\mathrm{d}x\right]+\\ +A\,\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})(x)\,\mathrm{d}x\right]\\ \times\int\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})(y)\,\mathrm{d}y\right]\,\mathrm{d}x,$ (77)
where $B$ is an integration constant.
$\bar{\eta}^{F}$ must be periodic. The first term on the right-hand-side of (77) is positive and periodic (it is the exponential of a periodic function). Notice that $f*\bar{\rho}$ is periodic, as a result of the circular convolution, and it sums to 0 when integrated over $\mathcal{S}$ due to Fubini’s theorem for convolutions. Hence, the integral of $f*\bar{\rho}$ is itself periodic. The second term on the right-hand-side of (77), cannot be periodic unless $A=0$, since $\mathrm{exp}\left\{\frac{1}{D}\int(f*\bar{\rho})(y)\,\mathrm{d}y\right\}$ is periodic, but it cannot sum to 0, being an exponential. Thus, $A=0$. For (77) to integrate to the non-plastic followers’ mass $\Phi^{F}$, we must have $B=\Phi_{F}/\int_{\mathcal{S}}h(x)\,\mathrm{d}x$, yielding the claim.∎

### v2

@@ 6 Higher-Dimensional Extension
Next, we extend the theoretical framework to periodic domains in higher-dimensions, $\Omega=[-\pi,\pi]^{d}$ ($d=2,3$). The model in (1) becomes
$\rho^{L}_{t}+\nabla\cdot[\rho^{L}\mathbf{u}+\rho^{L}(\mathbf{f}*\rho)]$ $=D\nabla^{2}\rho^{L}+q,$ (55a)
$\rho^{F}_{t}+\nabla\cdot[\rho^{F}(\mathbf{f}*\rho)]$ $=D\nabla^{2}\rho^{F}-q,$ (55b)
$\eta^{F}_{t}+\nabla\cdot[\eta^{F}(\mathbf{f}*\rho)]$ $=D\nabla^{2}\eta^{F},$ (55c)
@@ 6.1 Bio-inspired Control
$\nabla\cdot\left(\rho^{L}\mathbf{u}\right)=-Ke(\mathbf{x},t)-\nabla\cdot\left[\rho(\mathbf{f}*\rho)\right]{+}D\nabla^{2}\rho$ (56)
@@
The error dynamics obeys the higher-dimensional extension of (10). Choosing (56) transforms the error dynamics into (13), proving the claim. ∎
@@
To uniquely recover $\mathbf{u}$, a vector field, from the scalar relation (56), extra constraints need to be included, similar to [37]. Specifically, we set $\mathbf{w}:=\rho^{L}\mathbf{u}$, and
$Y=-Ke-\nabla\cdot\left[\rho(\mathbf{f}*\rho)(\mathbf{x},t)\right]{+}D\nabla^{2}\rho,$ (57)
@@
$\begin{cases}\nabla\cdot\mathbf{w}=Y,\\ \nabla\times\mathbf{w}=0,\end{cases}$ (58)
where we added a zero-curl condition to (56). Such a problem is analogous to the Poisson equation $\nabla^{2}\varphi=-Y$, where $\mathbf{w}=-\nabla\varphi$, and can be solved using Fourier series expansion. Once $\mathbf{\varphi}$ is recovered, $\mathbf{u}=-\nabla\varphi/\rho^{L}$.
@@
$q=\frac{1}{2}\,\nabla\cdot\left(\rho^{L}\mathbf{u}\right)+\frac{1}{2}\,\nabla\cdot\left[\rho^{*}(\mathbf{f}*\rho)\right]-\frac{D}{2}\nabla^{2}\rho^{*}+g,$ (59)
@@ 6.2 Stability Analysis
$\int(f_{1}*\psi)(x_{1},x_{2},x_{3})\,\mathrm{d}x_{1}=\int(f_{2}*\psi)(x_{1},x_{2}x_{3})\,\mathrm{d}x_{2}\\ =\int(f_{3}*\psi)(x_{1},x_{2},x_{3})\,\mathrm{d}x_{3},$ (60)
for any periodic $\psi$. Choosing $\mathbf{u}$ according to (56) (see Rem. 12) and $q$ as in (59) implies that $\bar{\rho}$ is a steady-state solution for the dynamics of $\rho$, with $\rho^{L},\rho^{F}>0$, and $\eta^{F}\geq 0$, if and only if the higher-dimensional extension of (20) holds.
@@
Under the additional assumption of isotropic kernel, the proof follows the same steps of those in Theorem 5. The only difference is the computation of the steady-state solution of $\eta^{F}$ (see (27)). Setting $\eta^{F}_{t}=0$ and $\rho=\bar{\rho}$ in (55c), we obtain
$\nabla\cdot\left[\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})\right]$ $=D\nabla^{2}\bar{\eta}^{F},$ (61)
@@
$\nabla\cdot\left[\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}\right]=0.$ (62)
Equation (62) is fulfilled if
$\nabla\bar{\eta}^{F}=({1}/{D})\,\bar{\eta}^{F}(\mathbf{f}*\bar{\rho}).$ (63)
@@
Many interaction kernels in the literature satisfy condition (60). Under this hypothesis, the steady-state solution $\bar{\eta}^{F}$ is uniquely defined, since $\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}=-\nabla\Psi$, where $\Psi$ is a harmonic scalar potential. This is analogous to couple (62) with $\nabla\times\left[\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}\right]=0$.
@@
the error dynamics globally converges to 0 in $\mathcal{L}^{2}(\Omega)$.
@@
The proof follows the same structure of that of Theorem 10. The time derivative of $V=\|e^{\eta}\|_{2}^{2}$ can be rewritten as
$V_{t}=-2D\|\nabla e^{\eta}\|_{2}^{2}-\int_{\Omega}\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\,\mathrm{d}\mathbf{x}\\ -2\mathrm{exp}(-Kt)\int_{\Omega}e^{\eta}\tilde{\mathbf{v}}\,\mathrm{d}\mathbf{x}{+\exp(-Kt)\int_{\Omega}(e^{\eta})^{2}\nabla\cdot\mathbf{v^{0}}\,\mathrm{d}\mathbf{x}},$ (68)
using the divergence theorem, vectorial identities, and posing $\mathbf{v}^{0}=\mathbf{f}*e^{0}$ and $\tilde{\mathbf{v}}=\nabla\cdot(\bar{\eta}^{F}\,\mathbf{v}^{0})$. The first term on the right-hand side can be bounded using the Poincaré-Wirtinger inequality, and for the second one the following bound holds:
$\left|\int_{\Omega}\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\,\mathrm{d}\mathbf{x}\right|\leq\int_{\Omega}\left|\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\right|\,\mathrm{d}\mathbf{x}\\ =\left\|e^{\eta}e^{\eta}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\right\|_{1}\leq\|e^{\eta}\|_{2}^{2}\|\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\|_{\infty}\\ \leq\|e^{\eta}\|_{2}^{2}\sum_{i=1}^{d}\|(f_{i}*\bar{\rho}_{x_{i}})\|_{\infty}\leq\|e^{\eta}\|_{2}^{2}\sum_{i=1}^{d}\|f_{i}\|_{2}\|\bar{\rho}_{x_{i}}\|_{2},$ (69)
@@
$\left|\exp(-Kt)\int_{\Omega}(e^{\eta})^{2}\nabla\cdot\mathbf{v}^{0}\,\mathrm{d}\mathbf{x}\right|\\ \leq\exp(-Kt)\|e^{\eta}\|_{2}^{2}\sum_{i=1}^{d}\|v_{i,x_{i}}^{0}\|_{\infty}$ (70)
$\left|-2\mathrm{exp}(-Kt)\int_{\Omega}e^{\eta}\tilde{\mathbf{v}}\,\mathrm{d}\mathbf{x}\right|\leq 2\mathrm{exp}(-Kt)\|e^{\eta}\tilde{\mathbf{v}}\|_{1}\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}\|_{2}\|\tilde{\mathbf{v}}\|_{2}$ (71)
@@
$V_{t}\leq\left(-2D+\sum_{i=1}^{d}\|f_{i}\|_{2}\|\bar{\rho}_{x_{i}}\|_{2}\right)V\\ {+\exp(-Kt)\sum_{i=1}^{d}\|v^{0}_{i,x_{i}}\|_{\infty}V}+2\mathrm{exp}(-Kt)\|\tilde{\mathbf{v}}\|_{2}\sqrt{V}.$ (72)
which is convergent due to Lemma 4 in [7].∎
@@ 6.3 Numerical validation
We consider a 2D mono-modal regulation scenario with a bivariate von Mises distribution (zero means, unit concentration). Agents interact through a 2D periodic Morse kernel ($L_{a}=\pi$, $L_{r}=\pi/4$, $\alpha=3.2$) with $D=0.05$, $\Phi^{F}=0.2$, $M^{L}(0)=M^{F}(0)=0.4$, $K=1$, $K_{FL}=1$, $K_{LF}=2$. Results in Fig. 4 for $t_{\mathrm{f}}=100$ are qualitatively comparable to those of 1D simulations in Figs 1. Convergence of $\rho$ to $\bar{\rho}$ occurs in  5 time units while convergence of $\rho^{L}$, $\rho^{F}$ and $\eta^{F}$ to their steady-state profiles is slower. We obtain final masses $M^{L}(t_{\mathrm{f}})\approx 0.26$ and $M^{F}(t_{\mathrm{f}})\approx 0.53$, consistent with the predicted $1/2$ ratio in Corollary 7.
@@ 7 Conclusions
We presented a bio-inspired leader-follower technique for spatial organization of large swarms via density control. Our strategy incorporates two crucial characteristics of natural systems: setting control objectives for the entire collective and introducing behavioral plasticity to tune the leaders-to-follower mass ratio. We derived conditions for existence and global stability of desired solutions, with numerical findings supporting the effectiveness, robustness, and versatility of our approach.
Our work does not come without limitations. Analytical convergence guarantees hold at the PDE scale (infinite swarms), and convergence from continuum to discrete agent-based implementations requires assessment. Theoretical results yield exact predictions only for steady-state mass ratios; numerical simulations suggest monotonic trends, but we cannot exclude cases with higher transient role-switching than predicted. Exploring these transient dynamics may inform energetic costs of behavioral plasticity. Furthermore, our control framework exploits full knowledge of the swarm parameters and dynamics, pointing at the necessity of a more robust architecture. Future work should address these limitations and extend continuum descriptions beyond spatial organization to complex tasks like self-assembly and collaborative manipulation [38].
@@ Appendix A Non-plastic followers at steady-state
$\bar{\eta}^{F}(x)=B\,\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})\,\mathrm{d}x\right]\\ +A\,\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})\,\mathrm{d}x\right]\int\mathrm{exp}\left[{-}\frac{1}{D}\int(f*\bar{\rho})\,\mathrm{d}y\right]\,\mathrm{d}x,$ (75)
where $B$ is an integration constant. $\bar{\eta}^{F}$ must be periodic. The first term on the right-hand-side of (75) is positive and periodic (it is the exponential of a periodic function). Notice that $f*\bar{\rho}$ is periodic, as a result of the circular convolution, and it sums to 0 when integrated over $\mathcal{S}$ due to Fubini’s theorem for convolutions. Hence, the integral of $f*\bar{\rho}$ is itself periodic. The second term on the right-hand-side of (75), cannot be periodic unless $A=0$, since $\mathrm{exp}\left\{{-}\frac{1}{D}\int(f*\bar{\rho})(y)\,\mathrm{d}y\right\}$ is periodic, but it cannot sum to 0, being an exponential. Thus, $A=0$. For (75) to integrate to the non-plastic followers’ mass $\Phi^{F}$, we must have $B=\Phi_{F}/\int_{\mathcal{S}}h(x)\,\mathrm{d}x$, yielding the claim.∎